OpenStudy (anonymous):

Solve –2x^2 – 16x – 44 = 0 x=-8+_i sqrt 6 x=-8+_ 2i sqrt 6 x=-4+_ i sqrt 6 x=-4+_ 2i sqrt 6

OpenStudy (anonymous):

–2x^2 – 16x – 44 = 0 x^2 + 8x + 22 = 0 (x + 4)^2 + 6 = 0 x + 4 = sqrt(-6) = plus minus i sqrt(6) x = -4 plus minus i sqrt(6)